OpenStudy (anonymous):

medal included!! Solve. s/39=18/26 s = ?

OpenStudy (anonymous):

@Kit_Kat_Nat_<3

OpenStudy (anonymous):

27

OpenStudy (anonymous):

s*26=39*18 s*26=702 s=702/26

OpenStudy (michele_laino):

@coolkilla03 please multiply both sides of your equality by 39

OpenStudy (anonymous):

39:26=3:2 13*3=39 13*2=26 S=9

OpenStudy (anonymous):

becuz the 39 is divided by 1.5 to get 26, so you flip it and use it on 18; 18 TIMES 1.5

OpenStudy (anonymous):

@Annie7077 is right... i didnt finish it sorry

OpenStudy (anonymous):

So s=27

OpenStudy (anonymous):

so the answer is 27

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):